### v2

@@ 5.2 Results
Grad. from program $j$ Grad. from top university
SPO SPO WPO WPO SPO SPO WPO WPO
UMAS UMAS UMAS UMAS
Total pairs 309 309 309 309 309 309 309 309
Bounds identify sign 206 142 26 10 101 77 7 3
Naive est. is outside the bounds 76 37 7 1 176 152 76 69
The table summarizes the results of the implementation of our method for 309 pairs $(j,k)$ and under different assumptions. For each of the two outcomes of interest and each of the four sets of assumptions, the table shows the numbers of pairs $(j,k)$ for which our estimated bounds identify the sign of the treatment effect of assignment to $j$ for those whose next-preferred option is $k$. It also shows the frequency at which naive RD estimates fall outside our bounds.
To further illustrate the informativeness and the relevance of our bounding approach, we look at 309 pairs $(j,k)$, which we select as follows. For each postsecondary program $j$ available on the platform at any point between 2004 and 2010 and each program $k$ ranked right after $j$ in a list, we compute the total number of students who listed $(j,k)$ in their ROLs between 2004 and 2010 and had a $j$-specific application score within a 30-point bandwidth of the year-specific cutoff for $j$. We find 327 pairs $(j,k)$ with at least 50 such applicants. These pairs involve 137 distinct programs $j$, each of them associated with a number of $k$s ranging from one to 18. We exclude the 18 pairs $(j,k)$ associated with one of these programs because no graduates are recorded in the data for this program. This yields a final set of 309 pairs involving 136 distinct $j$s.
As a simple procedure to construct a sample local to each admission cutoff $c_{j}$, we first use a 30-point bandwidth on either side of that cutoff. For a number of our $j$s of interest, there exists another program $\ell$ that uses the same placement score as $j$ and whose cutoff $c_{\ell}$ is within a 30-point distance of $c_{j}$. In these cases, we use a bandwidth smaller than 30 around the cutoff $c_{j}$ of interest, so it does not include any such $c_{\ell}$, and the bandwidth is possibly different across sides of $c_{j}$.
Table 2 provides statistics on the estimated bounds we obtain. The left panel shows statistics relative to our first outcome of interest, graduation from the treatment program $j$, while the right panel focuses on our second outcome of interest, graduation from a top university. Within each panel, statistics are shown for four different sets of assumptions: SPO with and without UMAS and WPO with and without UMAS. Under our strongest set of assumptions (SPO in combination with UMAS), our bounds identify the sign of the treatment effect on students’ probability of graduating from program $j$ (respectively, of graduating from a top university) for 206 (respectively, 101) out of the 309 pairs $(j,k)$ we study. This shows that our bounding approach is informative enough to recover the sign of causal effects for potentially many cases of interest. As we relax the assumptions we use to construct $\bm{\mathrm{Q}}_{j}$, the estimated bounds become less informative—they widen, and the frequency with which they recover the sign of the treatment effect decreases. This exercise can help empirical researchers evaluate and gauge the empirical content of the different assumptions they are willing to work with.
Finally, Table 2 also illustrates that treating behavior as naive rather than strategic can lead to misleading results.First, recall that the naive estimand $\mathbb{E}[Y\mid P_{j}=(j,k),S_{j}=c_{j}^{+}]-\mathbb{E}[Y\mid P_{j}=(j,k),S_{j}=c_{j}^{-}]$ identifies our causal parameter of interest if we assume everyone is a truth-teller. One might expect naive estimands to always lie within our bounds since truth-telling is a special case of our behavioral assumptions. However, Table 2 shows this is not true. Under the SPO and UMAS assumptions, the naive estimates fall outside our bounds for approximately 25% of the pairs of interest for our first outcome and for more than 50% of the pairs of interest for our second outcome. That the naive estimates fall outside our bounds provides additional evidence consistent with behavior being strategic. Indeed, our SD example in Section 4.1, right after Proposition 2, shows that the distribution of the potential outcomes conditional on reported local preferences and the priority score can change discontinuously at the cutoff in the presence of strategic behavior. In such a case, the naive estimand $\mathbb{E}[Y\mid P_{j}=(j,k),S_{j}=c_{j}^{+}]-\mathbb{E}[Y\mid P_{j}=(j,k),S_{j}=c_{j}^{-}]$ no longer equals the average treatment effect and may fall outside our bounds. More broadly, this also suggests that caution is warranted in cases where most but not all students submit $P$ with $|P|<K$ or very few students are assigned to their last-listed option. It may not be safe to simply assume that everyone is a truth-teller just because the cap does not bind for most people. In other words, naive estimates may be severely biased even if the share of constrained students is small.
@@ 6 Conclusion
This paper provides a novel approach to partially identifying the effects of mechanism assignment on future outcomes that is robust to strategic behavior. We illustrate our approach using data from a deferred-acceptance mechanism that assigns approximately 80,000 students to more than 1,000 university–major programs every year in Chile. First, we find substantial evidence of strategic behavior, confirming earlier findings from the literature on Chile. Second, we compute bounds on the average effect of program assignment on graduation outcomes. We do so for students whose scores are marginal to the cutoff for admission to a popular and selective program in the system: medicine at the Pontifical Catholic University in Santiago. Although admission to this program increases the likelihood of graduating from that same program, its effect on the likelihood of ever graduating from university varies with students’ second-best option. In this paper, we illustrate our bounding approach for RD-like parameters that are local to 309 university–major pairs. However, in many contexts, researchers may want to focus on more aggregate average effects, either because of lack of data near the multitude of cutoffs or because policymakers are primarily concerned with the returns to a change in the field of knowledge instead of a change in the university–major program, e.g., the impact of pursuing a STEM degree vs. a non-STEM degree. The RD identification strategy we propose in this paper is valid at the most granular level because this is the level at which students make decisions and the matching algorithm is run. The same strategy is not automatically justified for aggregates of university–major pairs. Aggregation of local effects brings additional challenges that we address in separate ongoing work.
Appendix A Proofs
A.1 Proof of Lemma 1

### v3

@@ 5.2.3 On the Importance of Accounting for Strategic Behavior
Graduation Graduation
from $j$ from top univ.
Total number of pairs $(j,k)$ 302 302
Number of cases with bounds identifying the TE sign 144 83
Number of cases with naive estimate outside the bounds 35 43
Share of cases with bounds identifying the TE sign 0.48 0.28
Share of cases with naive estimate outside the bounds 0.12 0.15
The table summarizes the results of the implementation of our method for 302 pairs $(j,k)$ under the SPO assumption. For each of the two outcomes of interest, the table shows the numbers of pairs $(j,k)$ for which our estimated bounds identify the sign of the treatment effect of assignment to $j$ for those whose next-preferred option is $k$. It also shows the frequency at which naive RD estimates fall outside our bounds.
@@ 6 Conclusion
This paper provides a novel approach to partially identify the effects of mechanism assignment on future outcomes that is robust to strategic behavior. We illustrate our approach using data from Chile, where a DA mechanism assigns about 80,000 students to more than 1,000 university-major programs every year, and for which, in line with previous literature, we find substantial evidence of strategic behavior. In two high-stakes settings—a selective entry program at the University of Chile and Medicine at PUC Santiago—we find heterogeneous effects of assignment on graduation outcomes across students’ preferred programs and next-best alternatives, consistent with preferences being linked to unobserved traits.
We illustrate our bounding approach for RD-like parameters at the granular university–program level, where students make choices and matching occurs. In many settings, however, researchers or policymakers may seek more aggregate effects—such as STEM vs. non-STEM—due to limited data near multiple cutoffs or broader policy relevance. Aggregation of local effects brings additional challenges that we address in separate ongoing work.
Supplemental Appendix for Online Publication
Causal Effects in Matching Mechanisms with Strategically Reported Preferences
By Marinho Bertanha, Margaux Luflade, and Ismael Mourifié.
Appendix A Additional Results Not in the Main Text
A.1 Parameter of Interest
This section complements and extends the discussion of our parameter of interest in Section 2 of the main text. The motivation behind our causal parameter is counterfactual analysis. Counterfactual analysis is a cornerstone of research in education market design, providing essential insights for policy evaluation and recommendation. In empirical applications, several types of counterfactual policies are commonly considered: changes to the assignment mechanism, modifications of priority scores, or adjustments in program capacities.
As detailed in Section V.B of Artemov et al. (2023), the literature offers two main approaches to counterfactual analysis. The first takes the submitted rank-ordered lists (ROLs), i.e., $P$, as inputs. This approach assumes that the submitted ROLs under the current policy reflect applicants’ true preferences and that individuals will submit the same lists under the counterfactual policy. However, Artemov et al. (2023) show the existence of robust equilibria in which a non-negligible fraction of participants do not submit their true preferences, even in a strategy-proof environment. [^17] Motivated by empirical evidence of non-truthful reporting that often has limited payoff consequences, Artemov et al. (2023) develop a novel solution concept—robust equilibrium—which relaxes the Bayesian Nash equilibrium by allowing for deviations from truthful reporting when the associated payoff losses are small. They argue that this finding raises important concerns about empirical methods that treat submitted ROLs as truthful representations of preferences—both under the current policy and in counterfactual scenarios.
The second approach, instead, assumes that individuals will report their true preference profile, $Q$ (potentially different from $P$), under the counterfactual policy. This approach requires the researcher to identify the true preference $Q$ in settings where agents behave strategically and misreport. Artemov et al. (2023) advocates for this second approach, arguing that a broad class of robust equilibrium concepts yield asymptotically stable matchings. The allocations produced by these equilibria can be well approximated by the stable matching induced under truthful reporting. In other words, this implies that for a wide range of reporting behaviors, the ultimate allocation will be stable, allowing us to consider counterfactual assignments as functions of the true preferences.
This paper examines the identification of parameters essential for counterfactual analyses following the second approach advocated in Artemov et al. (2023). Let  $(Y^{c},Q^{c},S^{c})$ denote the counterfactual distribution of outcomes, preferences, and scores in the population. [^18] For simplicity, we assume that $S_{j}=S$ for any $j$ in this section. Let  $\varphi^{c}(q,s)$ denote the program to which a student with true preference  $Q^{c}=q$ and grade  $S^{c}=s$ is assigned under the counterfactual assignment mechanism. [^19] This expression is derived under the assumption of a deterministic assignment mechanism, but it extends straightforwardly to cases involving extrinsic tie-breaking through randomization. For any function  $g\in\mathcal{G}$ of potential outcomes, we have:
$\mathbb{E}(g(Y^{c}))$ $=$ $\mathbb{E}\left(\mathbb{E}[g(Y^{c})|Q^{c},S^{c}]\right)$
$=$ $\sum_{j\in\mathcal{J}^{0}}\mathbb{E}(\mathbb{E}[\mathbb{I}\{\varphi^{c}(Q^{c},S^{c})=j\}g(Y^{c}(j))|Q^{c},S^{c}])$
$=$ $\sum_{j\in\mathcal{J}^{0}}\mathbb{E}(\mathbb{E}[\mathbb{I}\{\varphi^{c}(Q,S)=j\}g(Y(j))|Q,S])$
$=$ $\sum_{j\in\mathcal{J}^{0}}\mathbb{E}[w^{c}(j,Q,S)\mathbb{E}[g(Y(j))|Q,S]],$
where the weights  $w^{c}(j,q,s)=\mathbb{I}\{\varphi^{c}(q,s)=j\}$ are determined by the counterfactual assignment mechanism  $\varphi^{c}(q,s)$. The third equality holds under the so-called policy invariance assumption, i.e.,
$(Y^{c}(0),\ldots,Y^{c}(J),Q^{c},S^{c})\sim(Y(0),\ldots,Y(J),Q,S),$
which requires that the joint distribution of certain model primitives—such as potential outcomes, scores, and preferences—remains unchanged in the counterfactual world. See, for example, Heckman and Vytlacil (2005), who discuss both the necessity of this assumption and its limitations for certain classes of counterfactuals. Notice that, it may be more reasonable to assume that the true preference $Q$ is invariant to changes in the counterfactual world, i.e., $Q^{c}\sim Q$, rather than assuming that the submitted preferences $P$ remain unchanged. Indeed, some counterfactual scenarios may alter students’ strategic behavior, leading to $P^{c}\not\sim P$, which could pose challenges for researchers seeking to analyze such counterfactuals using the first approach described in Artemov et al. (2023).
Analyzing the welfare effects of a policy change involves comparing the counterfactual outcome distribution  $Y^{c}$ to the observed (status quo) outcome distribution  $Y$, through the quantity $\mathbb{E}[g(Y^{c})]-\mathbb{E}[g(Y)].$
As can be seen, the key ingredient needed to conduct this type of counterfactual analysis is knowledge of the average structural functions (ASF)  $s\mapsto\mathbb{E}[g(Y(j))\mid Q=q,S=s]$ for $j\in\{0,\ldots,J\}$ and $q\in\mathcal{Q}$. In practice, nonparametrically identifying each of these functions for all  $s\in\mathcal{S}$ is extremely challenging. However, an RD design can point or set identify differences of the ASF’s at a finite number of points, namely, $\mathbb{E}[g(Y(j))-g(Y(k))\mid Q=q,S=c_{j}]$ for certain values of $j$, $k$, and $q$. In data contexts with a large number of cutoffs, Bertanha (2020) proposes a consistent and asymptotically normal estimator for counterfactual weighted averages of these differences. Outside of these contexts, researchers may rely on parametric or semi-parametric assumptions to extrapolate information from finitely many points to the entire domain.
It is important to note that the continuity assumptions utilized by RD can point or set identify ASF’s for finitely many values of $j$ and $q$ but for infinitely many values of $s$ in certain subsets of $\mathcal{S}$. For example, in the SD example of the main text, $\mathbb{E}[g(Y(0))\mid Q=q,S=s]$ is identified for any $q$ and $s\leq c_{1}$. Researchers may then combine all the identifying information with assumptions on the ASF’s to construct bounds on $\mathbb{E}[g(Y^{c})]-\mathbb{E}[g(Y)].$ Following the ideas proposed by Mogstad et al. (2018), one may assume, for example, a Bernstein polynomial basis representation for each ASF:
$\mathbb{E}[g(Y(j))\mid Q=q,S=s]=\sum_{k=0}^{K}\theta_{jk}^{q}b_{k}(s),$
where  $\{b_{k}\}_{k=0}^{K}$ are known basis functions of $s$, and  $\theta_{j}^{q}:=(\theta_{j0}^{q},\ldots,\theta_{jK}^{q})$ are unknown parameters to be (set) identified.
Finally, due to data constraints inherent to the local nature of our RD estimation, our parameter of interest in the main text does not condition on the full preference profile  $Q=q$. Instead, we follow the intuition of Kirkeboen et al. (2016) and condition on what we define in Section 3 as the local preference  $Q_{j}$. One value of $Q_{j}$ maps to multiple values of $Q$ for individuals with $S=c_{j}$. It follows that our ASF parameters in the main text
$\mathbb{E}[g(Y(j))\mid Q_{j}=(j,k),S=c_{j}]~\text{ and }~\mathbb{E}[g(Y(k))\mid Q_{j}=(j,k),S=c_{j}]$
for $(j,k)\in\mathcal{P}$, equal weighted averages of $\mathbb{E}[g(Y(j))\mid Q,S=c_{j}]$ and $\mathbb{E}[g(Y(k))\mid Q,S=c_{j}]$ over values of $Q$ (see proof of Lemma 1 in Section B.1 of the appendix). Therefore, they contain identifying information about $\mathbb{E}[g(Y(j))\mid Q=q,S=s]$ and $\mathbb{E}[g(Y(k))\mid Q=q,S=s]$.
As discussed earlier, parameters of this form are essential for implementing a broad class of counterfactual analyses. Their identification is generally more challenging than that of parameters conditioned on $P$, such as $\mathbb{E}\left[g(Y(j))\,\middle|\,P_{j}=(j,k),\;S_{j}=c_{j}\right].$ We do not consider these parameters because they are not directly useful to the counterfactual approach advocated by Artemov et al. (2023) for the reasons given above.
A.2 Partial Orders Dominate Nonpartial Orders
$\varphi((p^{\prime},P^{-\omega_{0}}),\bm{\mathrm{S}})[\omega_{0}]~\bar{Q}(\omega_{0})~\varphi((p,P^{-\omega_{0}}),\bm{\mathrm{S}})[\omega_{0}]$ for every $P^{-\omega_{0}}$. In addition, we say $\varphi$ is strategy proof with unrestricted lists if submitting the true list of acceptable schools weakly dominates submitting anything else for every individual $\omega$.
Lemma 2.
Proof of Lemma 2:
This proof is from Haeringer and Klijn (2008), Lemma 4.2. We expand it here in terms of our framework and definitions of partial order.
Second, let $p^{\prime}$ be a weak partial order of $Q(\omega)$ that contains the schools listed in $\bar{p}$. Construct a new “true” preference ranking $q\subseteq\mathcal{J}^{0}$ as follows: (a) take the acceptable schools of $p^{\prime}$, and place them first in $q$ in the same order as they appear in $p^{\prime}$; (b) add a 0 to $q$ after the last school in part (a); (c) fill the remaining positions below 0 in $q$ with the schools not listed in $p^{\prime}$, in any order. Note that $p^{\prime}$ equals the true list of acceptable schools from $q$, $p^{\prime}$ is a weak partial order of $q$, and $p^{\prime}$ is the unique strong partial order of $q$.
The second claim of the lemma follows from the strategyproofness of $\varphi$ since the cap constraint is not binding for an individual whose number of acceptable schools is less than or equal to $K$ and submission of the true list of acceptable schools is feasible.
$\square$
A.3 Sharp Bounds on Treatment Effects
In this section, we utilize Artstein’s inequality (Theorem A.1 from Molinari (2020)) to characterize sharp bounds on the joint distribution of potential outcomes and true local preferences. That set of distributions produces sharp bounds on the averages of the treatment effects Y(j)Y(k)Y(j)-Y(k) conditional on $Q_{j}=(j,k)$ and $S_{j}=c_{j}$ for any comparable pair $(j,k)\in\mathcal{P}$. Recall that we denote the space of possible outcomes $Y$ as $\mathcal{Y}$. Let $2^{\mathcal{J}^{0}\times\mathcal{J}^{0}}$ denote the power set of $\mathcal{J}^{0}\times\mathcal{J}^{0}$} and $2^{\mathcal{Y}}$ denote the power set of $\mathcal{Y}$}.
Theorem 1.
Proof of Theorem 1: The random set $(\{Y\}\times\bm{\mathrm{Q}}_{j})$ is a measurable map from $\Omega$ to $\mathcal{Y}\times\mathcal{J}^{0}\times\mathcal{J}^{0}$, so it is compact valued. Following Definition A.1 from Molinari (2020), we say that $(\{Y\}\times\bm{\mathrm{Q}}_{j})$ is a random closed set because, for every compact set $\mathcal{K}\in\mathbb{R}^{3}$, the set $\{\omega\in\Omega:(\{Y(\omega)\}\times\bm{\mathrm{Q}}_{j}(\omega))\cap\mathcal{K}\neq\emptyset\}$ is a measurable event. By Assumption 6(i), the random vector $(Y,Q_{j})$ and the random set $(\{Y\}\times\bm{\mathrm{Q}}_{j})$ are measurable maps on the same probability space, and $\mathbb{P}\left[(Y,Q_{j})\in(\{Y\}\times\bm{\mathrm{Q}}_{j})|S_{j}=s\right]=1$ for any $s\in\mathcal{S}_{j}$.
Artstein’s inequality (Theorem A.1 from Molinari (2020)) characterizes the sharp set of all possible probability distributions for $(Y,Q_{j})$ that are consistent with our observation of $(\{Y\}\times\bm{\mathrm{Q}}_{j})$ and the fact that $\mathbb{P}[(Y,Q_{j})\in(\{Y\}\times\bm{\mathrm{Q}}_{j})|S_{j}=s]=1$. For any $s\in\mathcal{S}_{j}$, the inequality says that
$\mathbb{P}\left[Y\in A,\bm{\mathrm{Q}}_{j}\subseteq B|S_{j}=s\right]\leq\mathbb{P}\left[Y\in A,Q_{j}\in B|S_{j}=s\right]\;\;\;\forall A\in 2^{\mathcal{Y}},\;\ B\in 2^{\mathcal{J}^{0}\times\mathcal{J}^{0}}.$ (A-1)
Next, we use the cutoff characterization and rewrite the right-hand side of (A-1) as a sum. For $s\geq c_{j}$,
$=\sum_{(b,b^{\prime})\in B}\mathbb{P}\left[Y(b)\in A,Q_{j}=(b,b^{\prime})|S_{j}=s\right].$ (A-2)
Substitute (A-2) into (A-1) and take the limit as $s\downarrow c_{j}$ on both sides,
$\mathbb{P}\left[Y\in A,\bm{\mathrm{Q}}_{j}\subseteq B|S_{j}=c_{j}^{+}\right]\leq\sum_{(b,b^{\prime})\in B}\mathbb{P}\left[Y(b)\in A,Q_{j}=(b,b^{\prime})|S_{j}=c_{j}\right]\;\;\;\forall A\in 2^{\mathcal{Y}},\;\ B\in 2^{\mathcal{J}^{0}\times\mathcal{J}^{0}},$ (A-3)
Similarly, for $s>c_{j}$,
$=\sum_{(b,b^{\prime})\in B}\mathbb{P}\left[Y(b^{\prime})\in A,Q_{j}=(b,b^{\prime})|S_{j}=s\right].$ (A-4)
Use (A-4) into (A-1), and take the limit as $s\uparrow c_{j}$ on both sides,
$\mathbb{P}\left[Y\in A,\bm{\mathrm{Q}}_{j}\subseteq B|S_{j}=c_{j}^{-}\right]\leq\sum_{(b,b^{\prime})\in B}\mathbb{P}\left[Y(b^{\prime})\in A,Q_{j}=(b,b^{\prime})|S_{j}=c_{j}\right]\;\;\;\forall A\in 2^{\mathcal{Y}},\;\ B\in 2^{\mathcal{J}^{0}\times\mathcal{J}^{0}}.$ (A-5)
Lemma 1 of Chesher and Rosen (2017) applied to our case shows that the inequalities (A-3) and (A-5) are equivalent to:
$\square$
Theorem 1 characterizes the sharp set of all possible probability values of
$\mathbb{P}\left[Y(d)\in A,Q_{j}=(b,b^{\prime})|S_{j}=c_{j}\right]$ for $A\subseteq\mathcal{Y}$, $(b,b^{\prime})\in\mathcal{J}^{0}\times\mathcal{J}^{0}$, and $d\in\{b,b^{\prime}\}$. For a fixed $g\in\mathcal{G}$ of Assumption 2, that set of distributions allows us to define the sharp set of all possible means of potential outcomes,
$\mathbb{E}\left[g(Y(d))|Q_{j}=(j,k),S_{j}=c_{j}\right],$
$\mathbb{E}\left[g(Y(j))-g(Y(k))|Q_{j}=(j,k),S_{j}=c_{j}\right].$
In case of continuous $Y$, it is impossible to directly evaluate all inequalities of Theorem 1 because there are uncountably many sets $A\in 2^{\mathcal{Y}}$. This is one of the drawbacks of the Artstein inequality approach, which has been extensively discussed by Beresteanu et al. (2012). In case $Y$ takes finitely many values, e.g., when $Y$ is binary, the number of such inequalities is feasible to evaluate because $2^{\mathcal{Y}}\times\bm{\mathrm{\Lambda}}^{\cup+}_{j}$ (or $2^{\mathcal{Y}}\times\bm{\mathrm{\Lambda}}^{\cup-}_{j}$) has finitely many elements. In fact, the number of inequalities is slightly higher than the number of inequalities in Proposition 4 of the main text, which we utilize to compute lower bounds on $\delta_{j,k}^{+}$ and $\delta_{j,k}^{-}$.
A.4 Agents that Maximize Expected Utility
This section micro-founds the strategic reporting behavior of agents by describing in detail the problem they have to solve. We build on Agarwal and Somaini (2018) and Fack et al. (2019), who propose expected-utility maximizing agents that calculate probabilities of admission for each possible choice of rank-ordered list (ROL or $P$). In their settings, each agent knows her type, the mechanism, the distribution of other agents’ types, but does not know the individual types of other agents and the $P$’s they decide to submit. Thus, the uncertainty about an agent’s match comes from this lack of information about other agents.
Uncertainty about other people’s types and actions ultimately translates to uncertainty about the admission cutoffs ex-post the match. From an individual agent’s perspective, uncertainty about cutoffs is all that matters for the uncertainty of her final match. Therefore, we take a more “reduced-form” approach to the source of uncertainty in the agent’s problem and simply assume each agent has beliefs on ex-post cutoffs. This assumption not only simplifies the problem of the agent but is also more in line with the real-world decision process of students in Chile, who form expectations about cutoffs based on historical data and submit schools accordingly.
We now describe the problem of the agents (see Section A.5 for a numerical illustration). Before the matching mechanism is run, the agent knows her placement scores $\bm{\mathrm{s}}=(s_{1},\ldots,s_{J})$ and her true preferences $Q$ but does not know what the admission cutoffs will be after the matching is run. The agent sees the admission cutoffs as random variables $(C_{1},\ldots,C_{J})$. The strict preference relation $Q$ is represented by a vector of distinct utility values $(U_{0},U_{1},...,U_{J})$ so that $aQb~\Leftrightarrow~U_{a}>U_{b}$ for any $a,b\in\mathcal{J}^{0}$. We normalize $U_{0}=0$ for simplicity.
The agent has to decide on a ranking of acceptable schools $P$ to submit. The number of schools ranked in $P$ is $|P|$, and a feasible ranking has $1\leq|P|\leq K$. The set of all feasible rankings is defined as $\Delta P$. We let $P^{u}$ denote the $u$-th school listed in $P$, for $u=1,\ldots,|P|$. We define $L^{P}_{u}$ as the agent’s expected probability of being assigned to school $u$ when submitting ranking $P$, $u=1,\ldots,|P|$. $L^{P}_{0}$ denotes the expected probability of remaining unassigned, that is, of being matched to the outside option. Naturally, $L^{P}_{u}\geq 0$ for every $u$, and $\sum_{u=0}^{|P|}L^{P}_{u}=1$. Cutoff characterization wrt $P$ (Assumption 4) implies:
$\max\limits_{P\in\Delta P}\sum^{|P|}_{u=1}U_{P^{u}}L^{P}_{u}.$
A.5 Issues When Controlling for $P$
This section complements the discussion after Part III of the SD Example (Section 4.1) in the main text, where we point out the issue of controlling for $P_{4}=(4,2)$ in RD identification. We present an example of a continuum economy with agents that face cutoff uncertainty, maximize expected utility, and exhibit the type of discontinuous behavior that invalidates identification.
We build on the continuum economy of Section 2 and the agent optimization problem delineated in Section A.4. Part III of the SD Example (Section 4.1) considers an economy with $J=4$ schools and quota $K=3$ on submissions. Assume $S_{1}\sim U[0,1]$ and school capacities are $q_{1}=0.24$, $q_{2}=0.22$, $q_{3}=0.18$, and $q_{4}=0.16$. The outside option has unlimited capacity, i.e., $q_{0}=1$. The distribution of true preferences $Q$ is independent from $S_{1}$ and uniform over all permutations of $\{1,2,3,4\}$ with 0 in the last position (i.e., all schools are acceptable).
We first assume agents submit their true preferences and compute the four cutoffs by simulation. We draw $n=100,000$ agents from the population distribution of $(Q,S_{1})$ and run the SD algorithm. We obtain cutoffs $c_{1}=0.1988$, $c_{2}=0.2203$, $c_{3}=0.2958$, and $c_{4}=0.3623$. We call these “truth-telling cutoffs.”
True preferences can no longer be submitted in the presence of the quota $K=3$. Instead of listing all four acceptable schools, agents must think strategically and decide to drop at least one school from their lists. Agents have expectations about ex-post cutoffs and submit a partial order $P$ that maximizes their expected utilities (Section A.4). They submit a strong partial order of their true preferences, which implies that $|P|=3$. Expectations about ex-post cutoffs are modeled as random variables $C_{j}$, $j=1,\ldots,4$. We specify these distributions to be independent across $j$, each with a support containing the corresponding truth-telling cutoff $c_{j}$. All supports include a range of values so that no agent is certain about the ex-post cutoff. More specifically, there are two types of expectations and cardinal utility functions in this economy: those of agents with $Q=(4,3,2,1,0)$ and those of agents with $Q\neq(4,3,2,1,0)$.
Agents with $Q=(4,3,2,1,0)$ have expectations that are mixed continuous-discrete distributions. The discrete part of the distribution has one mass point at $c_{j}$ and the continuous part of the distribution is uniform over a closed interval. In particular, the mass points are $\mathbb{P}[C_{1}=c_{1}]=0.6$, $\mathbb{P}[C_{2}=c_{2}]=0.4$, $\mathbb{P}[C_{4}=c_{4}]=0.2$, while $C_{3}$ has no mass point; the continuous parts are uniform distributions $U[0;0.9]$ for $C_{1}$ and $C_{2}$, $U[0;1]$ for $C_{3}$, and $U[c_{4};0.85]$ for $C_{4}$. Their cardinal utilities are $4>3>2.99>2.98>0$, where $4$ is the cardinal utility of the first-ranked option in $Q$ and 0 is the cardinal utility of the last-ranked option in $Q$, that is, the outside option. All remaining agents, that is, those with $Q\neq(4,3,2,1,0)$, have expectations that are continuous distributions $C_{j}\sim U[c_{j}-0.15,c_{j}+0.15]$, for $j=1,2,3,4$. Their cardinal utilities are $4>3>2>1>0$, respectively, for the first-, second-, third-, fourth-, and fifth-ranked options in whatever $Q$ the agent has.
We compute optimal $P$s and ex-post cutoffs by simulation in the economy with strategic agents. First, we draw $n=100,000$ agents from the population distribution of $(Q,S_{1})$. Second, we compute the optimal preference submission $P$ of each agent drawn. Each individual chooses a strong partial order $P$ of her true preference ranking that maximizes her expected utility. The expected utility is the sum of the cardinal utilities weighted by assignment probabilities to each option listed in $P$. For an agent $i$ with score $s_{i}$ and cutoff expectations $C^{i}_{j}$, $j=1,2,3,4$, the probability of admission to the first-ranked option in a submission $P$ is $\mathbb{P}[s_{i}\geq C_{P^{1}}]$; the probability of admission into the second-ranked option is $\mathbb{P}[s_{i}<C_{P^{1}}]\mathbb{P}[s_{i}\geq C_{P^{2}}]$, etc, where $P^{u}$ denotes the $u$-th ranked option listed in $P$. Third and finally, we run the SD algorithm using $n$ observations of $(P,Q,S_{1})$. We obtain the same stable matching and cutoffs as in the truth-telling economy above.
Overall, utilities and expectations change with $Q$ but do not vary with $S_{1}$ once we condition on $Q$. By design, expectations and utilities of agents with $Q=(4,3,2,1,0)$ are different from the rest in order to have them behave in the discontinuous manner indicated in Part III of the SD Example. They optimally choose $P=(3,2,1)$ if $S_{1}<c_{4}$ but $P=(4,2,1)$ when $S_{1}\geq c_{4}$. Thus, if we control for $P_{4}=(4,2)$, the fraction of individuals with $Q_{4}=(4,3)$ jumps from zero to a positive number at the cutoff $c_{4}$, which invalidates the RD identification strategy. In fact, we find in the simulated economy
$\mathbb{P}[Q_{4}=(4,3)|S_{1}\in[c_{4},c_{4}+0.01],P_{4}=(4,2)]=0.3478$
$\mathbb{P}[Q_{4}=(4,3)|S_{1}\in[c_{4}-0.01,c_{4}),P_{4}=(4,2)]=0.$
There are two key ingredients that cause the discontinuous behavior of agents with $Q=(4,3,2,1,0)$. First, the mass points in their expectations indicate they possess a good amount of certainty about the ex-post values of the cutoffs of schools $1$, 22, and $4$. Second, their utility indicates that they are close to being indifferent between schools $1$, 22, and $3$. As the score increases and crosses the value $c_{4}$, the expected probability of having school $4$ in the budget set jumps from zero to 20%. The agent then decides to list school $4$ and rank it first, as it has the highest utility. There are three acceptable schools left for this agent, $3$, 22, and $1$, but only two spots left in her submission list given the quota $K=3$. The agent decides to skip school $3$ and list schools $2$ and $1$. The reason is twofold. First, the utility of $3$ is just a little higher than the utility of $2$ or $1$. Second, the probability of being assigned to either $2$ or $1$ when the agent submits $P=(4,2,1)$ is sufficiently higher than the probability of being assigned to either $3$ or $2$. This is the case because an agent with score $S_{1}=c_{4}$ is more certain to be above the cutoffs of schools $1$ or $2$ than to clear the cutoff of school $3$. Thus, the loss in expected utility of skipping school $3$ is more than compensated for by the gain of listing schools $2$ and $1$. Figure A-1 below illustrates these two points graphically.
Figure A-1: Expected Utilities and Probabilities of Assignment
Panel (a) depicts the expected utilities (y axis) from submitting four different strong partial orders as functions of the score $S_{1}$ (x axis) for an agent with $Q=(4,3,2,1,0)$. The dashed vertical lines represent the four cutoffs. The optimal submission is $P=(4,2,1)$ if $S_{1}\geq c_{4}$ and $P=(3,2,1)$ if $S_{1}<c_{4}$. For that same agent, panel (b) compares the probability of assignment (y axis) into schools $2$ or $1$ if the agent submits $P=(4,2,1)$ with the probability of assignment into schools $3$ or $2$ if the agent submits $P=(4,3,2)$. The probabilities are plotted as functions of the score $S_{1}$ (x axis) and the dashed vertical line represents cutoff $c_{4}$. These probabilities jump down at $c_{4}$ because the probability of being assigned to the first-best option $4$ jumps from zero to 20% at $c_{4}$ regardless if the agent submits $P=(4,2,1)$ or $P=(4,3,2)$. Panel (b) indicates that the agent expects a higher chance of being assigned to a school by submitting $P=(4,2,1)$ instead of $P=(4,3,2)$, thus the higher expected utility of $P=(4,2,1)$.
Appendix B Proofs of Results in the Main Text
B.1 Proof of Lemma 1